### v2

# Controlling noisy herds

Abstract
Controlling multi-agent systems is a persistent challenge in organismal, robotic, and social collectives, especially when agents exhibit stochastic indecisiveness – frequently switching between conflicting behavioral rules. Here, we investigate the control of such noisy indecisive collectives through the lens of century-old sheepdog trials, where small groups of sheep exhibit unpredictable switching between fleeing and following behaviors. Unlike cohesive large flocks, these small indecisive groups are difficult to control, yet skilled dog-handler teams excel at both herding and precisely splitting them (shedding) on demand. Using a stochastic model, we introduce two central parameters – pressure (stimulus intensity) and lightness (response isotropy) – to simulate and quantify herding and shedding dynamics. Light sheep rapidly reach stable herding states, while heavy sheep exhibit intermittent herding and orthogonal alignment to the dog. High response isotropy fosters group cohesion but complicates splitting tasks, highlighting the nuanced trade-offs in collective control of noisy herds. Surprisingly, we find that stochastic indecisiveness, typically perceived as a challenge, can be leveraged as a critical tool for efficient control, enabling controlled herding and splitting of noisy groups. Building on these insights, we develop the Indecisive Swarm Algorithm (ISA) for artificial agents and benchmark its performance against standard algorithms, including the Averaging-Based Swarm Algorithm (ASA) and the Leader-Follower Swarm Algorithm (LFSA). ISA minimizes control energy in trajectory-following tasks, outperforming alternatives under noisy conditions. By framing these results within a stochastic temporal network framework, we show that even with a probabilistic description of the future dynamics, network restructuring (temporality) enhances control efficiency in a specific class of control problems. These insights establish a scalable framework for controlling noisy, behavior-switching collectives, with applications in swarm robotics, cellular engineering, opinion dynamics, and temporal networks.
Control challenges in behavior-switching collectives
Emergent collective dynamics, where simple local interactions give rise to complex global behaviors, govern a wide range of systems. Examples include swarm robotics [1], animal collectives [2], social networks such as opinion dynamics [3], pedestrians’ movements [4], and vehicular traffic [5]. Controlling these systems poses significant challenges, as their behaviors often defy traditional control methods [6, 7, 8]. Unlike systems with predictable, linear dynamics, emergent systems require multiscale strategies that address both the microscopic interactions between individual agents and the macroscopic patterns that emerge at the group level [8].
Biological collectives exemplify these challenges and introduce additional complexity when individuals switch between behaviors unpredictably in response to environmental factors [9, 10, 11, 12, 13] (see SI Section 1 and Table S1 for a full list of behavior-switching systems from ants and locusts to seals and even humans). Schools of golden shiners (a type of fish) stochastically transition between swarming, schooling, and milling [14]; carrier ants transporting cargo alternate roles between lifters and pullers based on their orientation and the nest’s position [9], and sheep in small flocks randomly switch between leading and following roles [12]. These systems highlight the need for robust control strategies that account for the stochastic and context-dependent nature of individual behavior transitions and their cascading effects on group dynamics. Despite advances in control theory and swarm robotics [7, 8], managing these dynamics remains an open challenge, particularly in systems where behavioral switching transitions are inherently noisy.
Predator-prey systems provide a natural framework for studying the challenges of controlling noisy collectives with behavior-switching dynamics. For instance, flocks of starlings confuse raptors by transitioning between complex dynamical patterns. Similarly, large herds of wildebeests intermittently shift between selfish herding and solitary flight when confronted by predators like cheetahs. In response, predators adopt simplified strategies like focusing on a fixed point in space rather than tracking individual prey [15, 16] and split vulnerable individuals before leveraging speed and agility to secure their target [17, 18, 19]. These examples suggest that effective control of behavior-switching collectives does not always require precise tracking or prediction of behavioral transitions.
Unlike large, uncontrolled herds or purely theoretical systems, small flocks of sheep in competitions called sheep-dog trials provide a controlled environment where behavior-switching dynamics can be observed, quantified and systematically analyzed (see SI Section 2&3 for history and competition rules). At these trials, under threat, sheep oscillate stochastically between fleeing and regrouping, generating noisy, indecisive behaviors that challenge handlers and their dogs to cohesively herd or split groups under dynamic conditions (SI Video 1, 2). While prior studies have explored behavior-switching dynamics in the absence of a stimulus, in this work we uniquely frame sheep-dog trials as a control problem for "indecisive collectives" — systems where agents stochastically alternate between conflicting behaviors in presence of an external agent.
By introducing the metrics of "pressure" (stimulus intensity) and "lightness" (response isotropy), we bridge empirical observations with quantitative modeling to analyze various tasks in the sheepdog trials competition. Our findings reveal that indecision, often perceived as inefficiency, can paradoxically enhance control under specific conditions. Extending this framework, we propose a general Indecisive Swarm Algorithm (ISA) for artificial systems and demonstrate its effectiveness in trajectory-following tasks. By situating sheep-dog trials within the broader context of stochastic temporal networks, this work offers foundational principles for controlling noisy, behavior-switching collectives in applications ranging from swarm robotics to collective decision-making.
Sheep-dog Trials
Historically, shepherds have exploited predator-prey dynamics to control collectives, leveraging herding dogs to manage farm animals as early as 1700 BC (see Figure  1a, SI Section 2) [20]. When a solitary sheep encounters a threat, it flees; in large groups, sheep exhibit selfish herd behavior [21]. However, in small groups, sheep struggle to choose a survival strategy, indecisively switching between solitary and collective behaviors, creating unpredictability (Figure  1b and SI Video 1). This unpredictability inspired the creation of sheep-dog trials, a 100-year-old sport testing a dog’s ability to control small sheep groups ($N_{s}\leq 5$[22]. In these trials, handlers and dogs not only move sheep cohesively (called herding) (Figure  1c-e), but also split the groups into subgroups (called shedding) (Figure  1f-h), showcasing the dog’s skill in managing indecisive collectives (see SI Section 3)  [23].
Figure 1: Human-Dog-Sheep Interaction in Small Groups a A Bronze Age rock art panel at Valhaug on Jæren in southwestern Norway showing a shepherd herding a small group of sheep with the help of a dog (Photo Credit: Paul G. Keil) [23]. b Transition from single sheep response to large group response: While a single sheep flees under threat, sheep in a large group show selfish herd behavior. Sheep in small groups are highly indecisive and show a stochastic transition between the two behaviors, making the groups unpredictable. c-e Dynamics of herding in real sheep-dog system(SI video 3). f-h Dynamics of shedding in real sheep-dog system (SI video 3).
Building on this rich history, we systematically investigate the dynamics of sheep-dog interactions to understand how external control influences indecisive collectives. Sheep-dog trials offer a wealth of qualitative knowledge, developed over generations by dog trainers and shepherds, about these interactions. Terms such as "pressure" and "lightness" capture key aspects: pressure refers to the threat perceived by sheep from a dog’s actions, such as approaching, barking, or staring, while lightness describes the sheep’s responsiveness to these cues [24]. Trained dogs apply pressure to herd or shed (split) sheep, and handlers also contribute by exerting pressure through their body posture during shedding. Lightness is a measure of the responsiveness of the sheep. Light sheep respond to minimal pressure but may panic under high pressure, whereas heavy sheep resist until high pressure is applied directly from the front. Assessing sheep lightness early in trials is essential for achieving effective control [24].
Empirical Analysis of Orientation Dynamics to Inform Modeling
To examine how the control strategies of the dog differ between light and heavy sheep and to translate this nuanced qualitative knowledge into a quantitative framework, we recognize that herding and shedding both involve two sequential steps [25]. The first step, which we term the orientation step, involves nudging stationary sheep gently to induce directional change without causing panic. The second step, termed the movement step, involves increasing pressure to prompt movement (see SI section 4b and SI Video 2 for more details) [23]. Our study focuses only on the initial orientation step, isolating it from spatial dynamics such as movement and steric interactions, and considers only the orientation of sheep relative to the dog.
In sheep-dog trials, for a group of 5 sheep, a herding state is achieved when all the sheep are oriented away from the dog (Figure 1e). A shedding state occurs when the sheep divide into two groups of 3 and 2 individuals, with each group oriented perpendicularly away from both the dog and the handler (Figure 1h). To simplify our analysis, we classify sheep orientations into 4 directions relative to the dog: directly facing, perpendicular left, perpendicular right, and facing away (See SI Video 3, SI section 4a).
By quantifying the transitions between these 4 orientations in 21 videos of sheep-dog trials (See SI Section 4a for details), we observe clear differences in the behavior of light and heavy sheep during herding and shedding. In herding, light sheep quickly achieve a herding state and remain there, with occasional individual escapes. Heavy sheep, in contrast, exhibit intermittent herding states, often aligning orthogonally as a flock to the dog (Figure 3 a,c Bottom, SI Videos 6 and 7; Table S2). During shedding, light sheep frequently reorient individually, often due to panicking (Figure 3b Bottom). Heavy sheep, however, tend to align orthogonally to the dog and handler, synchronously switching between the two perpendicular directions away from the handler and the dog as a group (Figure 3b,d Bottom; SI Videos 8 and 9; Table S2).
The observed differences in behavior between light and heavy sheep, particularly their orientation transitions under varying pressure and lightness conditions, form the basis for defining the parameters in our stochastic model. These insights motivate the development of a quantitative framework to predict and generalize the indecisive behavior-switching dynamics observed in sheep-dog trials.
Figure 2: Quantitative Framework for Modeling the Orientation Step We simplify the model by making two assumptions: a. We only consider the orientation of the sheep and bin the 2D space in 4 directions, allowing us to model sheep as stationary pointers that can reorient in 4 possible directions. b. Transition rules describe how a sheep changes its direction when influenced by a dog/handler, other sheep, or spontaneously due to random noise with rates $\alpha_{ij},\gamma$, and $\epsilon$, respectively. The parameter $\alpha_{ij}$ represents the threat from the dog present in direction $j$ on the sheep oriented in direction $i$. When influenced by a dog/handler, sheep’s behavior changes depending on their orientation. Sheep facing the dog panic and randomly reorient, sheep perpendicular to the dog reorient to the opposite direction of the dog and sheep oriented away from the dog don’t change their orientation. c. Definition of lightness and responsiveness of sheep. Ideal light sheep with lightness $(L)=1$ respond to the dog irrespective of their orientation. Ideal heavy sheep with $L=0$ only respond if they are facing the dog. Sheep with intermediate lightness $10<L<$ 1 have higher responsiveness when facing the dog compared to being perpendicular to the dog d. Description of herding and shedding processes in our model. In herding, the goal is to align all the sheep away from the dog, whereas in shedding, the goal is to divide the group into two subgroups as required (typically into 3 and 2). Shedding involves both the handler and the dog.
Figure 3: Dynamics of Sheep Herding and Shedding. a-d Time-series of herding and shedding dynamics of a group of 5 sheep with $P=1$. Starting from a random initial orientation, we plot the time evolution of the number of sheep in each of the four directions. a-b herding and shedding of light sheep $L=0.9$. Top: Simulated time-series Bottom: Extracted transition time series from observed videos. Both in simulation and observations, once light sheep reach a herding state (all agents(sheep) in N direction (white)), they remain there with individuals escaping when influenced by noise. For shedding, light sheep panic and individually reorient without any discernible patterns. c,d Herding and shedding process for heavy sheep $L=0.1$. Top: Simulated time-series . Bottom: Extracted transition time series from observed videos. Heavy sheep show intermittent herding (all N) followed by synchronous alignment orthogonal to the dog (E or W) by all agents. For shedding, heavy sheep spend a significant time in E and W direction synchronously switching between them. The red vertical lines represent shedding events. Obtaining a shedding event in light sheep is difficult because they panic and randomly reorient when sandwiched between the handler and the dog. However, since heavy sheep synchronously switch between E and W, they provide narrow windows for the dog-handler teams to perform the shed. f,g Effect of pressure and lightness on $\tau_{stay}$ and $\tau_{reach}$ for herding and shedding, respectively. e,h Ease of herding ($E_{h}$) and ease of shedding ($E_{s}$) as functions of pressure and lightness. The result shows that it is easier to herd light sheep but easier to shed a group with an intermediate lightness ($L\approx 0.1$).
Modeling Indecisive Sheep Behavior
To generalize and predict the observed orientation dynamics, we next develop a stochastic model that formalizes the interplay of noise, social interactions, and external stimuli. This framework integrates qualitative insights from empirical observations with quantitative predictions for herding and shedding behaviors.
We model the indecisiveness in sheep behavior during the orientation step using a stochastic framework for $N_{s}$ stationary individual agents (sheep). These agents change orientations according to 3 rules:
  1. 1.
    Spontaneous Reorientation: Agents randomly change direction at a rate $\epsilon$ (noise).
  2. 2.
    Social Influence: Agents copy the orientation of neighboring agents at a rate $\gamma$.
  3. 3.
    External Stimulus Response: Agents reorient in response to external stimuli (dog or handler) at a rate $\alpha_{ik}$, where $i$ represents the agent’s current orientation, and $k$ represents the position of the stimulus (Figure 2b).
These reorientation rules are grounded in empirical observations of sheep-dog trials (see SI Video 10, SI Section 4c). For example, sheep facing the dog (frontal stimulus) panic and reorient randomly away from it, while those approached from the side reorient to face the direction opposite to the dog’s position. Conversely, sheep facing in the opposite direction from the stimulus remain unaffected (Figure 2b).
To simplify analysis and exploit rotational symmetry, without loss of generality, we fix the reference frame such that the external stimulus (dog) is always positioned in the South (S) direction as a convention. To compare our simulation with the experimental data, we bin the orientations of the simulated agents into 4 possible directions: North (N), South (S), East (E), and West (W) (Figure 2a). Unlike conventional Vicsek-type models, where agents average the influences from neighbors and stimuli at each time step (from now on referred to as averaging agents[26, 27], our approach allows agents to be influenced by one factor at a time, stochastically switching between them. This design builds on stochastic choice models that are widely used to capture stochastic switching dynamics in diverse systems across scales ranging from cancer cells and insects to fish schools and human opinion dynamics [28, 11, 29, 30, 31]. For sheep, this method incorporates prior evidence that small flock decision-making is Markovian, with frequent switches between leader and follower roles, even in the absence of external stimuli [32, 12].
Transition Dynamics and Governing Master Equation
The dynamics of sheep transitions between orientations can be described with the following reaction scheme:
$X_{i}\xrightarrow{\epsilon}X_{j\neq i}$ (1)
[…]
$X_{i}\xrightarrow{\alpha_{ii}}X_{j\neq i}\quad X_{i}\xrightarrow{\alpha_{i\tilde{i}}}\textit{No Effect}\quad X_{i}\xrightarrow{\alpha_{ik\neq{i,\tilde{i}}}}X_{\tilde{k}}.$ (3)
Here, $X_{i}$ represents a sheep in the $i^{\text{th}}$ direction, and $\alpha_{ik}$ is the influence of a stimulus located in direction $k$ on a sheep in orientation $X_{i}$. The notation $\tilde{i}$ indicates the direction opposite to $i$ (e.g., if $i=S$, then $\tilde{i}=N$). Importantly, the total number of sheep remains conserved, such that $\sum_{i}X_{i}=N_{s}$.
To capture the time evolution of these transitions, we model the system as a stochastic process governed by a master equation. The master equation is widely used across fields like reaction kinetics, population dynamics, and network theory to describe how probabilities of different system micro-states evolves over time [33, 34, 35, 36]. For sheep orientation dynamics, it captures the interplay of stochastic influences, social interactions, and external stimuli driving collective behaivor.
For this system, the master equation is expressed as:
$\frac{\partial}{\partial t}\mathbf{\mathcal{P}}(\bar{x},t)=\sum_{\bar{x}\neq\bar{x}^{\prime}}Tr(\bar{x}|\bar{x}^{\prime})\mathbf{\mathcal{P}}(\bar{x}^{\prime})-Tr(\bar{x}^{\prime}|\bar{x})\mathbf{\mathcal{P}}(\bar{x}).$ (4)
Here, $\bar{x}=\{x_{N},x_{S},x_{E},x_{W}\}$ represents the state of the system as the number distribution of sheep oriented in the four orientations, where $x_{i}$ denotes the number of sheep in the $i^{\text{th}}$ direction. $\mathbf{\mathcal{P}}(\bar{x},t)$ is the probability of observing the system in state $\bar{x}$ at time $t$ and $Tr(\bar{x}|\bar{x}^{\prime})$ is rate at which the system transitions from state $\bar{x}$ to $\bar{x}^{\prime}$.
The master equation captures the rate of change of the probability of observing the system in state $\bar{x}$ as the difference between processes bringing the systems into state $\bar{x}$ (first term in Equation (4)) and those moving the system away from it (second term in Equation (4)). This framework allows us to compute how the collective orientation of sheep evolves over time based on microscopic transition rates, including stochastic influences ($\epsilon$), social interactions ($\gamma$), and external stimuli ($\alpha_{ik}$). Detailed derivations of transition rates $Tr(\bar{x}|\bar{y})$ in terms of $\alpha_{ik},\gamma,\text{and}\epsilon$ are provided in SI Section 5.
Quantifying Pressure and Lightness Metrics
To bridge the microscopic transition rates of the master equations with the macroscopic behavior of the sheep, we introduce quantitative definitions for pressure and lightness:
  • Pressure ($P_{k}=\alpha_{kk}/\gamma$): Quantifies the relative influence of external stimuli (dog/handler) on a sheep facing the stimulus compared to the influence of neighboring sheep. Here, $\alpha_{kk}$ is the threat imposed by the stimulus on sheep facing it, while $\gamma$ denotes the influence of other sheep. Physically, $\alpha_{kk}$ can be interpreted as the threat imposed by a stimulus on a sheep looking towards it.
  • Lightness ($L=\alpha_{jk}/\alpha_{kk}$): Quantifies response isotropy, representing how a sheep oriented perpendicularly (E/W) responds to a stimulus compared to when it is directly facing it (S/N). Subscripts $j$ ad $k$ denote the sheep’s orientation and the stimulus’s position, respectively ($k=\{S,N\}$ for {dog, handler}).
Pressure $P_{k}$ ranges from 00 to $P_{max}$, where $P_{max}$ is the maximum pressure beyond which sheep begin to move. Since heavy sheep require a higher pressure to respond, $P_{max}$ depends on the sheep’s lightness. To compare the dynamics of light and heavy sheep and decouple pressure from lightness, we normalize $P$ by $P_{max}$.
Lightness $L$ ranges from 00 to $1$. For ideal light sheep $(L=1)$, the dog’s threat is isotropic and independent of orientation. In contrast, ideal heavy sheep $(L=0)$, only respond to stimuli from the front, ignoring stimuli from other directions (Figure 2c). Sheep with intermediate lightness values ($10<L<$ 1) reflect greater responsiveness to frontal stimuli compared to perpendicular stimuli (Figure 2b).
For simplicity, we assume a linear relationship between lightness $L$ and the influence of the dog $\alpha_{ik}$. Incorporating this allows us to consolidate transition rates as follows:
  • $\alpha_{ik}=\alpha$ if $i=k=\{S,N\}$
  • $\alpha_{ij}=L\alpha$ if $i=\{E,W\}$ and $j=\{S,N\}$
  • $\alpha_{ij}=0$ otherwise
Irrespective of the lightness, stimuli have no effect on sheep oriented opposite to them ($\alpha_{NS}=\alpha_{SN}=0$).
Dynamics of Herding and Shedding Sheep
We now use our stochastic model with 2 key parameters ($P$ $L$) and Gillespie’s algorithm [37] to simulate sheep dynamics using the master equation (4). Our analysis focuses on herding (orient all agents in $N$) and shedding (divide agents into $E$ $W$) behaviors for a small group size of $N_{s}=5$ under constant pressure. Despite its simplicity, the model predicts distinct behaviors for light and heavy sheep in both tasks.
In herding, light agents ($L=0.9$) quickly reach a herding state and remain stable, with occasional individual escapes driven by noise $\epsilon$ (Figure 3a Top). Heavy agents ($L=0.1$), in contrast, exhibit intermittent herding states and frequently align orthogonally to the dog (E or W) in a synchronous manner (Figure 3c Top). This behavior mirrors noise-induced switching observed in other small group systems (in the absence of external stimuli), where the reorientation of one individual triggers alignment changes across the group  [11, 29, 10]. Our model shows that isotropic responses facilitate herding, a result supported by empirical data from sheep-dog trials video (Figure 3a,c Bottom, SI Videos 6 and 7).
In shedding, light agents ($L=0.9$) frequently reorient without discernible patters due to their isotropic responsiveness (Figure 3b Top). Heavy agents ($L=0.1$), however, align orthogonally to the dog and handler and synchronously switch between E and W directions due to their selective responsiveness (Figure 3d Top), which closely matches the empirical shedding dynamics in sheep-dog trials (Figure 3b,d Bottom, SI Videos 8 and 9). Our model effectively captures nuanced behavioral differences between light and heavy sheep, during herding and shedding, demonstrating strong agreement with real-world observations (see SI Video 4 and Table S2).
To assess which sheep are easier to control, we quantify herding and shedding success by calculating reaching time ($\tau_{reach}$) and staying time ($\tau_{stay}$). Reaching time measures how long sheep take to achieve the desired orientation, while staying time indicates how long they remain in that state. Given the time-sensitive nature of sheep-dog trials, we define optimal conditions as those that maximize $\tau_{stay}$ and minimize $\tau_{reach}$. We calculate ease of herding (or shedding) as $E_{h(s)}=\tau_{\text{stay}}/\tau_{\text{reach}}$.
Our simulations reveal that in herding, increasing pressure or lightness reduces $\tau_{reach}$, while $\tau_{stay}$ depends only on noise $\epsilon$ (Figure 3f, and SI section 6). These results indicate that dogs use pressure to align sheep but cannot directly influence how long the alignment persists. The analysis of $E_{h}$ confirms that herding light sheep is easier than herding heavy sheep due to their uniform responsiveness to pressure (Figure 3e).
In shedding, the dog and handler create transient splits within the group, resulting in very short $\tau_{\text{stay}}$. Increasing pressure reduces $\tau_{reach}$, but higher lightness values lead to longer $\tau_{reach}$ (Figure 3g). The analysis of $E_{s}$ shows that shedding very heavy or very light sheep is particularly challenging (Figure 3h). In trials, shedding tasks push the capabilities of the dog-handler team to their limits, as the dog counteracts the sheep’s selfish herding tendencies. The model also predicts that the optimal pressure for both herding and shedding is the maximum stationary pressure $P_{\text{max}}$. Beyond this threshold, sheep flee uncontrollably (Figure 3e,h). In practice, dogs dynamically adjust pressure to account for sheep heterogeneity and changes in lightness, underscoring the complexity of controlling indecisive collectives.
Figure 4: Comparison of averaging and indecisive sheep with two-step control: a Herding performance is compared under low noise conditions ($\epsilon/\gamma=0.08$, shown in magenta, top) and high noise conditions ($\epsilon/\gamma=0.8$, shown in teal, bottom). Starting from random initial orientations, both averaging and indecisive agents reach the target under low noise, with averaging agents doing so more quickly. However, under high noise, averaging agents fail to reach the target due to corruption by noise, while indecisive agents, though slower, successfully reach the target. b In shedding tasks, averaging agents fail to split the group, whereas, indecisive agents consistently succeed (SI Video 5) example in figure and SI Video 5: $\epsilon/\gamma=0.8$ and $\alpha/\gamma=1$).
Figure 5: Comparison of indecisive and averaging sheep a,b Schematic representations of averaging and indecisive sheep agents. Averaging agents average the influence of all factors to update their orientation, while indecisive agents stochastically switch between single influencing factors. c-d Time series showing the dynamics of a small flock of indecisive and averaging sheep during herding and shedding processes for $\alpha=\gamma=\epsilon=0.1$. e,f Evaluation of ease of herding ($E_{h}$) and ease of shedding ($E_{s}$) for both. Averaging sheep are easier to herd under low relative noise, but indecisive sheep are easier to herd in high-noise conditions. For shedding, indecisiveness is crucial since averaging agents fail to split. The teal and magenta squares in e represent the low noise and high noise herding dynamics described in Figure 4a
Is Indecisiveness Useful?
Does indecisiveness only pose challenges, or can it aid the dog in controlling the flock? To investigate this, we next simulate both the orientation and movement steps of indecisive sheep dynamics in a 2D arena. We extend the 4-direction stochastic framework into continuous 2D space, enabling agents to asynchronously update their orientation to any direction between $-\pi$ to $\pi$, rather than limiting them to discrete directions (N, S, E, W). The rules from Figure 2b remain unchanged, but when sheep panic and reorient due to random noise ($\epsilon$) or face the stimulus $\alpha_{kk}$, they randomly select a new direction within the range $-\pi$ to $\pi$, excluding their current orientation.
We compare these indecisive agents with standard averaging agents (Vicsek-type model), where sheep agents synchronously update their position by averaging the effects of all influencing factors (Figure 5a,b) [38, 39, 26, 40]. Both models incorporate alignment with other sheep, repulsion from the dog, and random noise. For consistency, we focus on ideal light sheep ($L=1$).
We first simulate the herding problem with the two models (Figure 4a). Due to asynchronous updates, indecisive sheep agents move in random directions and diverge, rendering the flock uncontrollable (SI video 5). This result emphasizes the necessity of the two-step control process implemented by shepherd dogs. Integrating this insight into our simulations replicates real dog-herding behaviors, demonstrating that effective control of noisy, indecisive collectives like sheep flocks requires independent regulation of movement and orientation (SI Video 5).
To compare the controllability of the two models, we simulate the herding process under different noise levels: $\epsilon/\gamma=0.08$ (Figure 4a, top) and $\epsilon/\gamma=0.8$ (Figure 4a, bottom). Starting from a random initial orientation, we find that at low relative noise, averaging agents outperform indecisive agents, reaching the target faster (Figure 4a, “initial" and “final"). However, at high relative noise, averaging agents fail to reach the target, while indecisive agents, although slower, successfully complete the task. In shedding tasks, averaging agents fail to split in two subgroups under all noise levels, whereas, indecisive agents consistently succeed, regardless of the magnitude of relative noise (Figure 4b).
To quantify control efficiency, we use the metrics ease of herding (shedding), $E_{h(s)}=\tau_{stay}/\tau_{reach}$. Here, $\tau_{reach}$ is the time required for the flock to achieve the desired orientation (all agents aligned away from the dog in herding, or split into 3 and 2 agents away from the dog and handler in shedding), and $\tau_{stay}$ is the duration the flock maintains the preferred orientation.
By analyzing the time series of agent orientations, we observe that in herding, averaging agents maintain orientation but are corrupted by noise, while indecisive agents alternate between epochs of perfect herding and random reorientation (Figure 5c). Using $E_{h}$, we identify which model is easier to control under varying pressure ($P=\alpha/\gamma$) and relative noise ($\epsilon/\gamma$). We observe a phase transition: averaging agents are easier to herd at low noise, but indecisive agents outperform them as noise increases (Figure 5e). For shedding, regardless of pressure or noise, the averaging agents fail to split, whereas indecisive agents consistently split (Figure 5d,f). These results demonstrate that introducing indecisiveness improves control for complex tasks involving both herding and splitting, particularly under noisy conditions.
Exploiting Indecisiveness for Better Control
Having demonstrated that indecisiveness can aid control, we extend this concept to artificial systems for broader applications. Specifically, we investigate whether stochastic indecisiveness can improve control strategies in robotics, particularly for multi-agent systems navigating constrained trajectories. To test this, we consider the classic trajectory-following problem, a widely studied challenge in robotics [41, 42, 43]. We model a swarm of agents without visual sensing (blind agents) that rely solely on communicated orientations from other agents. A controller agent with visual sensing capability is tasked to steer the swarm from an initial position to a target position along a predefined trajectory by applying repulsion forces on individual agents.
To evaluate the role of indecisiveness, we develop the Indecisive Swarm Algorithm (ISA), where agents stochastically switch inputs between either the controller or another agent Figure 6a. ISA agents exhibit two key differences from noisy sheep: (1) they update their dynamics synchronously, and (2) they do not panic in response to the controller or randomly change orientation ($\epsilon=0$), hence can be programmed to deterministically move away from the controller. Thus, stochastically switching input sources (controller/agent) remains the sole source of randomness for agents. We benchmark ISA against two standard algorithms: a Vicsek-type Averaging Swarm Algorithm (ASA), where agents average inputs from all sources [44, 26], and the Leader-Follower Swarm Algorithm (LFSA), where agents form a fixed hierarchical network, either copying another agent or following the controller non-reciprocally [45, 46, 47].
All three algorithms share two key parameters; i.e., repulsion from the controller $\tilde{\alpha}$ and alignment with another agent $\tilde{\gamma}$. The dynamics of the three algorithms are expressed as:
$\theta_{ASA}^{(i)}(t)=-\tilde{\alpha}\theta_{ac}+\sum_{agents}\tilde{\gamma}\theta_{aa}$ (5)
$\theta_{ISA}^{(i)}(t)=\begin{cases}-\theta_{ac}&\text{with probability }\tilde{\alpha},\\ \theta_{aa}&\text{with probability }\tilde{\gamma},\\ \end{cases}$ (6)
$\theta_{LFSA}^{(i)}(t)=\begin{cases}-\theta_{ac}&\text{if }a^{(i)}\text{ is leader},\\ \theta_{aa}&\text{if }a^{(i)}\text{ is follower},\\ \end{cases}$ (7)
where
$a^{(i)}=\begin{cases}\text{leader}&\text{with probability }\tilde{\alpha},\\ \text{follower}&\text{with probability }\tilde{\gamma}.\\ \end{cases}$
Here $\theta_{alg.}^{(i)}(t)$ represents the orientation of agent $a_{i}$ at time $t$ under each algorithm. $\theta_{ac}$ and $\theta_{aa}$ denote the orientation of the agent due to repulsion from the controller and alignment with other agents, respectively. The normalized parameters $\tilde{\alpha}$ and $\tilde{\gamma}$ are the weights with which agents respond to the controller and align with others, respectively such that $\tilde{\alpha}+\tilde{\gamma}=1$. We define the stimulus intensity $I=\tilde{\alpha}/\tilde{\gamma}$, which generalizes the pressure $(P)$ used in the sheep-dog trials.
Figure 6b compares the trajectories of ISA, ASA, and LFSA for a group size $N_{s}=50$. At high stimulus intensity ($I$), all algorithms guide agents along the pre-defined trajectory effectively. However, at low stimulus intensity, ASA and LFSA agents deviate significantly, while ISA agents remain on track, demonstrating the utility of indecisiveness in reducing control effort (SI Video 11).
Temporality and Control Energy
To better understand why ISA performs better than ASA/LFSA, we examine the algorithms as special cases of stochastic non-reciprocal temporal networks [48], where interactions between agents (nodes) evolve dynamically. Such networks are characterized by two timescales: the timescale at which the network restructures $(\tau_{n})$ and the timescale at which agents update their dynamics $(\tau_{d})$. Agents update their dynamics every $(\tau_{d})$ by averaging interactions between consecutive updates (Figure 6a). To capture the relationship between these timescales, we define temporality $(\mathcal{T})$ as:
$\mathcal{T}=\tau_{d}/\tau_{n}$ (8)
This temporality parameter enables us to interpolate between the behaviours of different swarm algorithms (Figure 6a). When $\mathcal{T}\rightarrow 0$ ($\tau_{n}>>\tau_{d}$), the system mimics LFSA, where agents interact non-reciprocally in a fixed network topology. When $\mathcal{T}\rightarrow\infty$ ($\tau_{n}<<\tau_{d}$), agents average all inputs over time, resembling ASA. At $\mathcal{T}=1$ ($\tau_{n}=\tau_{d}$), where network restructuring and dynamics updates synchronize, ISA emerges as a distinct behaviour.
Since we have framed the problem within a temporal network framework, we can apply the concept of control energy, a measure widely used to assess the controllability of complex networks [49, 48]. To calculate the control energy required to herd a swarm of agents, we define a safe path as the area between two boundary lines similar to ref. [41]. The controller’s task is to move the swarm to a target position while keeping the swarm’s center of mass within the safe path. To prevent divergence, we also impose bounds on the variance of the swarm (Figure  6c-Top).
At each dynamics update ($\tau_{d}$), the controller begins with a low stimulus intensity $I$ and systematically increases $I$ until the swarm moves along the constrained trajectory. For a swarm that reaches the target successfully, we calculate the control energy $\mathcal{E}$ [48] as:
$\mathcal{E}=\sum_{t}\frac{1}{2}I_{min}(t)^{2}$ (9)
where $I_{min}(t)$ represents the minimum stimulus intensity required at time $t$ for the swarm to follow the path. The swarm trajectories and the corresponding variation of $I_{min}(t)$ over time are shown in Figure  6c (top and bottom, respectively).
For LFSA, fixed pairwise non-reciprocal interactions often cause the swarm to split into small clusters, preventing it from reaching the target (SI video 11). ASA agents, on the other hand, exhibit strong averaging behavior (strong ordering), which makes them prone to overshooting and alignment challenges at the trajectory boundaries. This requires the controller to apply high input intensities to reorient the agents in the preferred direction, resulting in peaks in $I_{min}(t)$ ISA agents, however, reduce the likelihood of cluster formation through network restructuring while maintaining sufficient flexibility to avoid the strong order of ASA agents. As a result, ISA agents successfully reach the target with significantly lower $I_{min}(t)$ compared to ASA.
To evaluate if ISA is the optimal strategy for controlling noisy swarms, we calculate the control energy $\mathcal{E}$ as a function of temporality $\mathcal{T}$ and group size $N_{s}$. Figure 6d shows the variation of $\mathcal{E}$ and the fraction of failed trajectories across different $\mathcal{T}$ values. For $N_{s}=50$, the failure fraction increases as $\mathcal{T}$ decreases, reflecting the limitations of LFSA. Regardless of group size, the control energy reaches its minimum at $\mathcal{T}=1$, highlighting the optimality of ISA for herding noisy swarms along predefined paths.
Figure 6: Indecisive Swarm Algorithm: a A schematic illustrates how temporality $\mathcal{T}$ influences a swarm of agents steered by a controller. Each snapshot depicts the agents’ orientation, the controller’s relative position, and their interactions. The network restructures every $\tau_{n}$, and agents update their directions at every dynamic update $\tau_{d}$ by considering interactions occurring between consecutive dynamic updates. When $\mathcal{T}=3$, the network restructures three times between two dynamics updates, and agents calculate a weighted average of all interactions at each dynamic update to determine their new direction. As $\mathcal{T}\rightarrow\infty$, this behavior resembles an averaging swarm algorithm (ASA), where agents average inputs from all the others and the controller’s repulsion to update their direction. In contrast, when $\mathcal{T}=0.5$, the network updates every two dynamic updates. As $\mathcal{T}\rightarrow 0$, the system mimics a leader-follower swarm algorithm (LFSA), where agents randomly follow a chosen agent or respond to the controller, indefinitely. At $\mathcal{T}=1$, the system operates as an indecisive swarm algorithm (ISA), with each network update directly followed by a dynamic update. b Trajectories of LFSA, ISA, and ASA agents are shown for different stimulus intensities $I$ with $N_{s}=50$. At high stimulus intensity, all agents follow the predefined path. At low stimulus intensity, only ISA agents stay on the predefined path, while ASA and LFSA agents deviate significantly. c-Top. The constrained predefined path used for control energy $\mathcal{E}$ calculations. ISA and ASA agents successfully reach the target, while LFSA agents fail. c-Bottom $I_{min}(t)$ for ISA and ASA agents. The peak in $I_{min}$ for ASA corresponds to the moment when the swarm reaches the boundary. d $\mathcal{E}$ and failure fraction shown as a functions of $\mathcal{T}$ for 1,500 simulations. As $\mathcal{T}<1$, the failure fraction increases for large group sizes ($N_{s}=50$). Control energy $\mathcal{E}$ achieves a minimum at $\mathcal{T}=1$, demonstrating the optimality of ISA for herding noisy agents along predefined trajectories.
Discussion
Summary
This study investigates control mechanisms for noisy, indecisive collectives, using sheepdog trials as a model system. These trials challenge trained shepherd dogs to herd and shed (split) small flocks of sheep ($N_{s}\leq 5$), where the dynamics differ markedly from larger flocks. Unlike the cohesive selfish herd behavior seen in large groups, sheep in small flocks stochastically transition between fleeing (solitary behavior) and following the group (collective behavior), making them harder to control (i.e., an indecisive herd). By combining qualitative insights from expert dog handlers with a stochastic modeling framework, we analyze how trained dogs manage these indecisive sheep collectives.
We find that sheep behavior depends on two key factors: the dog’s threat level and the sheep’s switching dynamics. Within the shepherding community, these factors are encapsulated by the terms “pressure" (the dog’s threat) and “lightness" (the isotropy of the sheep’s responsiveness). Light and heavy sheep exhibit distinct behaviors during herding and shedding tasks. To translate this nuanced qualitative knowledge into a quantitative framework, we developed a stochastic model to describe indecisive sheep behavior. The model reveals that trained dogs employ a two-step control strategy: first aligning stationary sheep to a desired orientation (orientation step) before increasing threat to initiate movement (movement step). Focusing on the orientation step, we modeled sheep as stationary agents that reorient stochastically. This analysis formalized the concepts of pressure and lightness, confirming their utility as core descriptors of sheep behavior. Comparing the model to data from actual sheepdog trials, we find that high isotropy aids group cohesion (for herding) but complicates splitting, while the dynamics of indecisive sheep are largely governed by the two parameters, pressure and lightness.
We also investigated whether indecisiveness benefits the controller rather than solely posing a challenge. Extending our framework to simulate both orientation and movement steps in a 2D arena, we compared indecisive sheep agents with standard averaging-based Vicsek-type agents. While averaging agents outperform indecisive agents under low noise conditions, the reverse is true at higher noise levels. For shedding tasks, averaging agents consistently fail to split, while indecisive agents shed easily, irrespective of noise levels. These results highlight how trained dogs exploit the sheep indecisiveness as a tool and underscore the importance of the two-step control process.
Finally, we explored whether indecisiveness could improve control strategies in artificial systems. Developing the Indecisive Swarm Algorithm (ISA), we compared it against the Averaging-based Swarm Algorithm (ASA) and Leader-Follower Swarm Algorithm (LFSA) in a trajectory-following task. ISA agents successfully followed predefined trajectories at low stimulus intensities from the controller, unlike ASA and LFSA agents, which deviated significantly. Framing swarm algorithms as stochastic temporal networks, we identified two tunable timescales: the dynamics update timescale $(\tau_{d})$ and network restructuring timescale $(\tau_{n})$. By defining temporality $\mathcal{T}=\tau_{d}/\tau_{n}$, we showed that adjusting $\mathcal{T}$ reproduces all three algorithms: ASA ($\mathcal{T}\to\infty$), LFSA ($\mathcal{T}\to 0$), and ISA ($\mathcal{T}=1$). Borrowing the concept of control energy from control theory, we quantified the stimulus intensity required to steer a swarm. ISA required the least control energy, demonstrating its effectiveness in herding noisy swarms.
Our findings reveal the counterintuitive advantages of indecisiveness in controlling noisy collectives, with applications ranging from sheepdog trials to artificial swarms. By introducing deliberate indecisiveness, controllers can enhance their ability to perform complex tasks, such as herding and splitting, while also reducing effort in simpler tasks like trajectory-following.
Why Sheepdog Trials are Challenging
If indecisive agents require less control effort, why are sheepdog trials considered so challenging? To address this, we extended our indecisive model to large group sizes (SI Section 8). While the model was originally designed to explain the behavior of small groups ($N_{s}\leq 5$) in response to external stimuli, its extension to larger group sizes captures dynamics consistent with known sheep behaviors. This broader application allowed us to propose a unified phase diagram for indecisive behavior (see SI Section 8 for details), offering insights into transitions between different behavioral regimes as group size and stimulus specificity change.
The phase diagram (Figure 7) illustrates the likelihood of individuals being influenced by controlling stimuli ($\alpha$), intra-group interactions ($\gamma$), or random noise (non-specific stimulus) ($\epsilon$). Stimulus specificity, defined as the ratio of $\alpha/\epsilon$, measures the strength of external stimuli to noise. External stimuli, such as a dog’s pressure or the departure of an informed sheep, are key factors driving transitions between behaviors.
We identify three distinct behavioral regimes: flocking (red), dominated by intra-group interaction, resulting in cohesive group behavior; fleeing (blue), dominated by specific stimuli where individuals act independently, ignoring the group; and grazing (green), dominated by random noise, with individuals disregarding both specific stimuli and the group.
In small groups, increasing stimulus specificity shifts behavior from grazing to fleeing. In larger groups, flocking dominates under typical stimulus intensities. However, when stimulus specificity becomes extremely high - such as during a predator attack or an encounter with an untrained dog - the flocking phase transitions to fleeing, even in large groups (Figure 7).
We validated our model’s predictions by comparing them with prior empirical studies of sheep behavior. King et al. [21] (circle) observed that intermediate-sized groups (46 sheep) exhibited selfish herd behavior under high stimulus specificity, with herding dogs inducing cohesion. Toulet et al. [50] (square) found that when a trained sheep departs intermediate-sized groups (8-32 sheep), the group reaches a consensus to follow or ignore the individual, demonstrating the dominance of intra-group interactions even under mild stimuli (low specificity). Ginelli et al. [51] and Gomez-Nava et al. [12] (star and triangle) studied group dynamics without external stimuli. Ginelli focused on large groups (100 sheep), while Gomez-Nava examined small groups (4 sheep). Both identifyied intermittent grazing and flocking epochs, aligning with the grazing-flocking transition boundary in our model. These behaviors suggest an evolutionary anticipation of external threats as a defense mechanism.
Our model (red line) predicts that small groups transition from grazing to uncontrolled fleeing through a narrow flocking phase as external stimulus increases. This prediction explains why managing small flocks is particularly difficult in sheepdog trials. Since individual sheep vary in their responsiveness to stimuli, effectively herding or splitting small flocks requires the dog to balance intra-group cohesion with individual responsiveness, as excessive stimulus risks triggering chaotic fleeing. This underscores the complexity of controlling small, indecisive collectives, where behavioral transitions depend on a delicate interplay of external stimuli, noise, and group interactions.
Figure 7: Unifying Phase Diagram of Indecisive Collective Behavior. We present a qualitative phase diagram for indecisive collective behavior as a function of group size ($N_{s}$) and specificity of external stimulus ($\alpha/\epsilon$) and compare it with the behavior of sheep presented in previous works. We demonstrate that even if indecisive collective model doesn’t explicitly explain the selfish herd behavior of sheep in large group sizes, an extension of the model to incorporate group sizes across scales show show behavior matching the behavior of sheep reported in the literature. Three distinct regimes - fleeing (blue, $\alpha$ dominated), flocking (blue, $\gamma$ dominated), and grazing (green, $\epsilon$ dominated) - are shown. Black shapes represent literature results for different $N_{s}$ and stimulus specificity ($\alpha/\epsilon$). The $\bullet$ indicates King et. al.’s [21] finding that intermediate groups (46 sheep) exhibit selfish herd behavior under threat. The $\blacksquare$ represents Toulet et. al.’s [50] study of consensus in intermediate (8 to 32 sheep) following a trained sheep. The $\bigstar$ and $\blacktriangle$ denote Ginelli et. al.’s [51] and Gomez-Nava et. al.’s [12] findings of intermittent flocking and grazing epochs in large (100 sheep) and small groups (4 sheep), respectively. The red line shows the behavior range in sheepdog trials, transitioning from grazing to uncontrolled fleeing through a narrow flocking phase, highlighting the difficulty of managing small indecisive groups of sheep.
Temporality and Indecisiveness
Temporal networks have been shown to require significantly less control energy than static networks [48]. This efficiency arises from their ability to leverage changing topologies to exploit favorable configurations, thereby reducing the need to counteract unfavorable system dynamics. In contrast, static networks, with their fixed structures, often force controllers to expend substantial energy to navigate energetically costly directions or to overcome inherent system dynamics. A useful analogy is sailing, where adjusting the sail to align with shifting wind directions enhances efficiency, rather than struggling against them [48].
However, this framework assumes that the controller has prior knowledge of future topology changes. Without such foresight, temporality can actually increase control energy by orders of magnitude compared to static networks [52]. This raises a key question: can temporality still offer advantages in the absence of knowledge about future changes?
We demonstrate that for a specific class of control problems – herding of indecisive collectives – temporality can significantly reduce control energy, even without prior knowledge of topology changes. Herding differs from standard control problems in complex networks in two fundamental ways. First, while traditional controllability in the context of complex networks involves steering the system from any initial state to any desired state within the state space [8, 53], herding focuses on guiding all agents (nodes) to a fixed consensus state along a predefined trajectory. Second, unlike standard control, where the controller can choose which nodes receive input [8], in herding, the controller has no choice over which node receives the control input.
Our analysis reveals that indecisive collective – stochastic temporal networks with restructuring timescales equal to system dynamics timescales ($\mathcal{T}=1$) – are optimal for minimizing control energy. This finding offers a new perspective on leveraging temporality for efficient control of noisy living and robotic swarms, even in the absence of topology foresight.
Broader Implications and Future Directions
Without external stimuli (e.g., a dog or a handler), our indecisive model extends a general stochastic framework widely applied across diverse systems, including auto-catalytic biochemical reactions [54], heterogeneous cancer cell populations [28], collective animal movement [29, 10, 11], and human opinion dynamics [55] (SI Section 1, Table S1). By introducing the concept of an external controller, or "shepherd," our analysis establishes a foundational framework for controlling noisy groups in a variety of domains. For instance, Zajdel et al.[56] demonstrated a shepherd-dog-inspired mechanism to guide cells along specified trajectories, highlighting the potential for shepherding strategies in cellular systems. Building on these insights, our framework could guide the design of effective control mechanisms to herd and sort heterogeneous cell collectives. Such strategies hold promise for applications like promoting wound healing through coordinated cell movement or selectively isolating healthy cells from infected populations. More broadly, our approach bridges seemingly disparate fields, providing a foundation for algorithms capable of effectively controlling stochastic, indecisive swarms.
While we presented a simplified model to explore the effects of sheep indecisiveness in sheep-dog-handler interactions, the real-world dynamics of this system are far more intricate. Shepherd dogs can instinctively predict sheep movements, but expertly trained dogs uniquely integrate instinct with handler commands to achieve precise coordination. In successful teams, the handler and dog operate cohesively, eliminating the need for constant monitoring. Instead, they function as a unified entity, sharing cognitive resources to analyze and anticipate the sheep’s behavior in real time [23, 57, 24]. Systematically studying these sophisticated interactions – spanning verbal, physical, and visual modalities – could inspire advances in swarms robotics. Embedding distributed cognitive mechanisms across diverse agents, including drones, quadruped robots, and human or AI operators, could enable more adaptive, robust, and noise-tolerant systems capable of handling internal and external uncertainties effectively [58, 59].
Acknowledgements
We thank the Australian National Sheepdog Trials for allowing us to use the data from their YouTube channel. We also thank Ceri S. Rundle for allowing us to use the data of the sheepdog trials "A way with the dogs" on her YouTube channel CSJ Specialist Canine Feeds. We express our gratitude to Doyle Ivie of Woodsend Stock Dogs for sharing his expertise on sheepdog trials and dog training. We also thank Matthew Bull, Benjamin Seleb, Atanu Chatterjee, Pankaj Rohilla, and Ishant Tiwari for the discussions and feedback on the manuscript. S.B. acknowledges funding support from NIH MIRA Grant R35GM142588; NSF Grants PHY-2310691; CMMI-2218382; CAREER iOS-1941933; and the Open Philanthropy Project.
Author contributions
T.C. and S.B. conceptualized the study. T.C. derived the mathematical models, performed the simulations, and analyzed the data. Both the authors contributed to writing the manuscript. S.B. supervised the project.
Competing interests
The authors declare no competing interests.
Materials & Correspondence
The codes are available at https://github.com/bhamla-lab/Controlling-Noisy-Herds-2025

### v3

@@
-Controlling noisy herds
+Temporal network restructuring improves control of indecisive collectives
 Abstract
-Controlling multi-agent systems is a persistent challenge in organismal, robotic, and social collectives, especially when agents exhibit stochastic indecisiveness – frequently switching between conflicting behavioral rules.
-Here, we investigate the control of such noisy indecisive collectives through the lens of century-old sheepdog trials, where small groups of sheep exhibit unpredictable switching between fleeing and following behaviors.
-Unlike cohesive large flocks, these small indecisive groups are difficult to control, yet skilled dog-handler teams excel at both herding and precisely splitting them (shedding) on demand.
-Using a stochastic model, we introduce two central parameters – pressure (stimulus intensity) and lightness (response isotropy) – to simulate and quantify herding and shedding dynamics.
-Light sheep rapidly reach stable herding states, while heavy sheep exhibit intermittent herding and orthogonal alignment to the dog.
-High response isotropy fosters group cohesion but complicates splitting tasks, highlighting the nuanced trade-offs in collective control of noisy herds.
-Surprisingly, we find that stochastic indecisiveness, typically perceived as a challenge, can be leveraged as a critical tool for efficient control, enabling controlled herding and splitting of noisy groups.
-Building on these insights, we develop the Indecisive Swarm Algorithm (ISA) for artificial agents and benchmark its performance against standard algorithms, including the Averaging-Based Swarm Algorithm (ASA) and the Leader-Follower Swarm Algorithm (LFSA).
-ISA minimizes control energy in trajectory-following tasks, outperforming alternatives under noisy conditions.
-By framing these results within a stochastic temporal network framework, we show that even with a probabilistic description of the future dynamics, network restructuring (temporality) enhances control efficiency in a specific class of control problems.
-These insights establish a scalable framework for controlling noisy, behavior-switching collectives, with applications in swarm robotics, cellular engineering, opinion dynamics, and temporal networks.
-Control challenges in behavior-switching collectives
+Controlling stochastic temporal networks remains an open challenge in control theory.
+While predictable temporal networks with known future dynamics enhance controllability, real-world networks often exhibit stochasticity and unpredictability, making control harder.
+Here, we investigate control mechanisms for stochastic temporal networks by analyzing how biological controllers, such as shepherd dogs, manage panicked flocks of sheep.
+We studied a century-old shepherding competition, the sheepdog trials, where small groups of sheep unpredictably switch between fleeing and following behaviors–effectively forming stochastic temporal networks.
+Unlike large, cohesive flocks, these small, indecisive flocks are difficult to control, yet skilled dog-handler teams excel at both herding and splitting them (shedding) on demand.
+Using a stochastic choice model to describe the sheep’s behavioral shifts, we found that trained dogs exploit stochastic indecisiveness, typically seen as an obstacle, as a control tool, enabling both herding and splitting of noisy groups of sheep.
+Building on these insights, we developed the Indecisive Swarm Algorithm (ISA) for artificial agents and benchmarked its performance against standard approaches, including the Averaging-Based Swarm Algorithm (ASA) and the Leader-Follower Swarm Algorithm (LFSA).
+ISA minimizes control energy in trajectory-following tasks and outperforms alternatives under noisy conditions.
+By framing these results within a stochastic temporal network perspective, we demonstrate that even probabilistic knowledge of future dynamics can enhance control efficiency in specific scenarios.
+These findings establish a framework for managing stochastic temporal networks with applications in noisy, behavior-switching animal collectives, swarm robotics, and opinion dynamics.
+Introduction
 Emergent collective dynamics, where simple local interactions give rise to complex global behaviors, govern a wide range of systems.
 Examples include swarm robotics [1], animal collectives [2], social networks such as opinion dynamics [3], pedestrians’ movements [4], and vehicular traffic [5].
-Controlling these systems poses significant challenges, as their behaviors often defy traditional control methods [6, 7, 8].
-Unlike systems with predictable, linear dynamics, emergent systems require multiscale strategies that address both the microscopic interactions between individual agents and the macroscopic patterns that emerge at the group level [8].
-Biological collectives exemplify these challenges and introduce additional complexity when individuals switch between behaviors unpredictably in response to environmental factors [9, 10, 11, 12, 13] (see SI Section 1 and Table S1 for a full list of behavior-switching systems from ants and locusts to seals and even humans).
-Schools of golden shiners (a type of fish) stochastically transition between swarming, schooling, and milling [14]; carrier ants transporting cargo alternate roles between lifters and pullers based on their orientation and the nest’s position [9], and sheep in small flocks randomly switch between leading and following roles [12].
-These systems highlight the need for robust control strategies that account for the stochastic and context-dependent nature of individual behavior transitions and their cascading effects on group dynamics.
-Despite advances in control theory and swarm robotics [7, 8], managing these dynamics remains an open challenge, particularly in systems where behavioral switching transitions are inherently noisy.
-Predator-prey systems provide a natural framework for studying the challenges of controlling noisy collectives with behavior-switching dynamics.
-For instance, flocks of starlings confuse raptors by transitioning between complex dynamical patterns.
+Controlling these systems is challenging, as their behaviors often defy traditional control methods [6, 7, 8].
+Unlike systems with predictable, linear dynamics, emergent systems are best described as complex networks that require multiscale strategies to address both the microscopic interactions between individual agents and the macroscopic patterns that emerge at the group level [8].
+Most of these networked systems introduce additional complexity when individual agents (nodes) switch between different behaviors, leading to temporal restructuring in the network.
+Biological collectives and social interactions in humans serve as prime examples of such behavior switching  [9, 10, 11, 12, 13] (see SI Section 1 and Table S1 for a full list of behavior-switching systems from ants and locusts to seals and humans).
+Carrier ants transporting cargo alternate roles between lifters and pullers based on their orientation and the nest’s position [9], sheep in small flocks randomly switch between leading and following roles [12], and during an epidemic outbreak, humans frequently switch between different interaction partners, facilitating spread of diseases [14].
+These systems highlight the need for control strategies that account for the stochastic and context-dependent nature of individual behavior transitions and their cascading effects on the evolving temporal networks, where edges dynamically reorganize over time [15, 16, 17].
+Recently, it has been shown that temporal restructuring can improve the controllability of a network [18].
+Specifically, temporal networks require less time and less energetic cost to be controlled than their static counterparts [19, 20].
+This counterintuitive observation relies on the fact that the future dynamics of the network are predictable and are exploited in designing the controls in the previous steps.
+However, when the switching dynamics are stochastic and unpredictable, temporality can make the control process more energetically demanding compared to a static network [21].
+Therefore, despite advances in control theory and swarm robotics [7, 8], managing the dynamics of stochastic temporal networks remains an open challenge, particularly in systems where individual agents exhibit behavior switching.
+Predator-prey systems provide a natural framework for studying the challenges of controlling such noisy networks with behavior-switching dynamics.
+For instance, flocks of starlings confuse raptors by transitioning between complex dynamic patterns.
 Similarly, large herds of wildebeests intermittently shift between selfish herding and solitary flight when confronted by predators like cheetahs.
-In response, predators adopt simplified strategies like focusing on a fixed point in space rather than tracking individual prey [15, 16] and split vulnerable individuals before leveraging speed and agility to secure their target [17, 18, 19].
-These examples suggest that effective control of behavior-switching collectives does not always require precise tracking or prediction of behavioral transitions.
-Unlike large, uncontrolled herds or purely theoretical systems, small flocks of sheep in competitions called sheep-dog trials provide a controlled environment where behavior-switching dynamics can be observed, quantified and systematically analyzed (see SI Section 2&3 for history and competition rules).
-At these trials, under threat, sheep oscillate stochastically between fleeing and regrouping, generating noisy, indecisive behaviors that challenge handlers and their dogs to cohesively herd or split groups under dynamic conditions (SI Video 1, 2).
-While prior studies have explored behavior-switching dynamics in the absence of a stimulus, in this work we uniquely frame sheep-dog trials as a control problem for "indecisive collectives" — systems where agents stochastically alternate between conflicting behaviors in presence of an external agent.
-By introducing the metrics of "pressure" (stimulus intensity) and "lightness" (response isotropy), we bridge empirical observations with quantitative modeling to analyze various tasks in the sheepdog trials competition.
-Our findings reveal that indecision, often perceived as inefficiency, can paradoxically enhance control under specific conditions.
-Extending this framework, we propose a general Indecisive Swarm Algorithm (ISA) for artificial systems and demonstrate its effectiveness in trajectory-following tasks.
-By situating sheep-dog trials within the broader context of stochastic temporal networks, this work offers foundational principles for controlling noisy, behavior-switching collectives in applications ranging from swarm robotics to collective decision-making.
-Sheep-dog Trials
+In response, predators, instead of complex control mechanisms, adopt simplified strategies like focusing on a fixed point in space rather than tracking individual prey [22, 23].
+This allows them to split vulnerable individuals before leveraging speed and agility to secure their target [24, 25, 26].
+These examples suggest that effective control of stochastic temporal networks with behavior-switching individuals does not always require precise prediction of behavioral transitions.
+In this work, we analyze such control mechanisms by studying shepherd dogs managing small flocks of sheep in a competition called the sheep-dog trials.
+Two key features of these competitions make them model systems for investigating control mechanisms in stochastic temporal networks.
+First, during these trials, when threatened, panicked sheep oscillate unpredictably and indecisively between fleeing from the dog and following other sheep, forming a stochastic temporal network.
+Trained shepherd dogs are highly effective at managing these noisy flocks under fluctuating conditions (SI Video 1).
+Second, unlike interactions between predators and large herds of animals in the wild, the sheepdog trials competitions provide a controlled environment where the behavior-switching dynamics of the sheep can be observed, quantified, and analyzed (see SI Section 2&3 for history and competition rules).
+By bridging empirical observations with quantitative modeling to analyze various tasks in the sheepdog trials competition, we find that shepherd dogs utilize the behavior switching in sheep for herding and splitting (shedding) the flocks.
+Behavior switching dynamics have been previously studied in the context of animal collectives and human societies (voter models) using individual-based stochastic choice models [27, 11, 28, 29, 30].
+In this work, we build on the existing framework to frame sheep-dog trials as a control problem for "indecisive collectives" — systems where agents stochastically alternate between different behaviors and interaction partners in the presence of an external agent.
+This paper is structured as follows: We begin by exploring the nuances and rules of sheepdog trials.
+Next, we present a stochastic framework to develop quantitative metrics such as "pressure" and "lightness" that capture the nuanced behavior of sheep.
+The framework is based on qualitative insights from experienced handlers, and empirical data on sheep-dog dynamics.
+We then present a stochastic choice model and the master equation to model the indecisive transitions in sheep movement, comparing our model’s predictions with observed dynamics.
+Building on this, we investigate whether sheep indecisiveness could benefit the dog.
+Our findings reveal that stochastic indecisiveness can aid the dog in both herding and shedding tasks.
+Finally, we extend our analysis to develop the Indecisive Swarm Algorithm (ISA), a swarm control strategy inspired by shepherding dynamics.
+By modeling ISA as a non-reciprocal stochastic temporal network and comparing it against the standard Averaging-Based Algorithm (ASA) and Leader-Follower Swarm Algorithm (LFSA), we demonstrate that for specific control tasks like herding, ISA minimizes control energy requirements.
+Results
+Terminology in Sheep-dog Trials
 Historically, shepherds have exploited predator-prey dynamics to control collectives, leveraging herding dogs to manage farm animals as early as 1700 BC (see Figure  1a, SI Section 2) [31].
@@
 However, in small groups, sheep struggle to choose a survival strategy, indecisively switching between solitary and collective behaviors, creating unpredictability (Figure  1b and SI Video 1).
-This unpredictability inspired the creation of sheep-dog trials, a 100-year-old sport testing a dog’s ability to control small sheep groups ($N_{s}\leq 5$[22].
-In these trials, handlers and dogs not only move sheep cohesively (called herding) (Figure  1c-e), but also split the groups into subgroups (called shedding) (Figure  1f-h), showcasing the dog’s skill in managing indecisive collectives (see SI Section 3)  [23].
+This unpredictability led to the creation of sheep-dog trials, a 100-year-old sporting competition testing a dog’s ability to control small sheep groups ($N_{s}\leq 5$[33].
+In these trials, handlers and dogs not only move sheep cohesively (called herding), but also split the groups into subgroups (called shedding), showcasing the dog’s skill in managing indecisive collectives (Figure  1c-h, see SI Section 3)  [34].
 Figure 1: Human-Dog-Sheep Interaction in Small Groups a A Bronze Age rock art panel at Valhaug on Jæren in southwestern Norway showing a shepherd herding a small group of sheep with the help of a dog (Photo Credit: Paul G.
@@
 f-h Dynamics of shedding in real sheep-dog system (SI video 3).
-Building on this rich history, we systematically investigate the dynamics of sheep-dog interactions to understand how external control influences indecisive collectives.
-Sheep-dog trials offer a wealth of qualitative knowledge, developed over generations by dog trainers and shepherds, about these interactions.
-Terms such as "pressure" and "lightness" capture key aspects: pressure refers to the threat perceived by sheep from a dog’s actions, such as approaching, barking, or staring, while lightness describes the sheep’s responsiveness to these cues [24].
+In sheep-dog trials, qualitative terms such as "pressure" and "lightness" convey the following aspects: pressure refers to the threat perceived by sheep from a dog’s actions, such as approaching, barking, or staring, while lightness describes the sheep’s responsiveness to these cues [35].
 Trained dogs apply pressure to herd or shed (split) sheep, and handlers also contribute by exerting pressure through their body posture during shedding.
@@
 To compare our simulation with the experimental data, we bin the orientations of the simulated agents into 4 possible directions: North (N), South (S), East (E), and West (W) (Figure 2a).
-Unlike conventional Vicsek-type models, where agents average the influences from neighbors and stimuli at each time step (from now on referred to as averaging agents[26, 27], our approach allows agents to be influenced by one factor at a time, stochastically switching between them.
-This design builds on stochastic choice models that are widely used to capture stochastic switching dynamics in diverse systems across scales ranging from cancer cells and insects to fish schools and human opinion dynamics [28, 11, 29, 30, 31].
+Traditionally, the dynamics of sheep are modeled with a conventional averaging-based approach, where each sheep averages the orientations of its neighbors and threat from the dog before deciding the direction of motion at each time step (from now on referred to as averaging agents)  [37, 38].
+While this fully connected static network approach captures the dynamics of large flocks of sheep, it fails to simulate the indecisive behavior-switching dynamics in small flocks.
+Our approach allows agents to be influenced by one factor at a time, stochastically switching between them.
+This design is an extension of stochastic choice models that are widely used to capture stochastic switching dynamics in diverse systems across scales ranging from cancer cells and insects to fish schools and human opinion dynamics (such as two state voter models) [27, 11, 28, 29, 30].
 For sheep, this method incorporates prior evidence that small flock decision-making is Markovian, with frequent switches between leader and follower roles, even in the absence of external stimuli [39, 12].
@@
 To capture the time evolution of these transitions, we model the system as a stochastic process governed by a master equation.
-The master equation is widely used across fields like reaction kinetics, population dynamics, and network theory to describe how probabilities of different system micro-states evolves over time [33, 34, 35, 36].
+The master equation is widely used across fields like reaction kinetics, population dynamics, and network theory to describe how probabilities of different system micro-states evolve over time [40, 41, 42, 43].
 For sheep orientation dynamics, it captures the interplay of stochastic influences, social interactions, and external stimuli driving collective behaivor.
@@
 The teal and magenta squares in e represent the low noise and high noise herding dynamics described in Figure 4a
-Is Indecisiveness Useful?
+Can Indecisiveness Improve Control?
 Does indecisiveness only pose challenges, or can it aid the dog in controlling the flock?
@@
 However, at high relative noise, averaging agents fail to reach the target, while indecisive agents, although slower, successfully complete the task.
-In shedding tasks, averaging agents fail to split in two subgroups under all noise levels, whereas, indecisive agents consistently succeed, regardless of the magnitude of relative noise (Figure 4b).
+In shedding tasks, averaging agents fail to split into two subgroups under all noise levels, whereas, indecisive agents consistently succeed, regardless of the magnitude of relative noise (Figure 4b).
 To quantify control efficiency, we use the metrics ease of herding (shedding), $E_{h(s)}=\tau_{stay}/\tau_{reach}$.
@@
 These results demonstrate that introducing indecisiveness improves control for complex tasks involving both herding and splitting, particularly under noisy conditions.
-Exploiting Indecisiveness for Better Control
-Having demonstrated that indecisiveness can aid control, we extend this concept to artificial systems for broader applications.
-Specifically, we investigate whether stochastic indecisiveness can improve control strategies in robotics, particularly for multi-agent systems navigating constrained trajectories.
-To test this, we consider the classic trajectory-following problem, a widely studied challenge in robotics [41, 42, 43].
+Developing an Indecisive Swarm Algorithm (ISA)
+Having demonstrated that indecisiveness can improve control, we extend this concept to artificial systems for broader applications.
+Specifically, we investigate whether stochastic indecisiveness can improve control strategies in robotics, particularly for multi-agent networks navigating constrained trajectories.
+Control mechanisms that are effective for single agents often fail in multi-agent networked systems due to the emergence of complex global behaviors from simple local interactions between agents.
+Designing controllers capable of simultaneously managing local, microscopic interactions and global network-level dynamics is still an open problem in control theory [8, 48].
+We show that by introducing stochastic indecisiveness in the multi-agent system, one can enable simple controllers, designed to control single agents, to control multi-agent networked systems.
+To illustrate this, we use the classic trajectory-following problem, a widely studied challenge in robotics [49, 50, 51].
 We model a swarm of agents without visual sensing (blind agents) that rely solely on communicated orientations from other agents.
 A controller agent with visual sensing capability is tasked to steer the swarm from an initial position to a target position along a predefined trajectory by applying repulsion forces on individual agents.
+The problem is a special class of controllability problems called herdability [52, 53, 54].
 To evaluate the role of indecisiveness, we develop the Indecisive Swarm Algorithm (ISA), where agents stochastically switch inputs between either the controller or another agent Figure 6a.
@@
 The swarm trajectories and the corresponding variation of $I_{min}(t)$ over time are shown in Figure  6c (top and bottom, respectively).
-For LFSA, fixed pairwise non-reciprocal interactions often cause the swarm to split into small clusters, preventing it from reaching the target (SI video 11).
-ASA agents, on the other hand, exhibit strong averaging behavior (strong ordering), which makes them prone to overshooting and alignment challenges at the trajectory boundaries.
-This requires the controller to apply high input intensities to reorient the agents in the preferred direction, resulting in peaks in $I_{min}(t)$ ISA agents, however, reduce the likelihood of cluster formation through network restructuring while maintaining sufficient flexibility to avoid the strong order of ASA agents.
-As a result, ISA agents successfully reach the target with significantly lower $I_{min}(t)$ compared to ASA.
 To evaluate if ISA is the optimal strategy for controlling noisy swarms, we calculate the control energy $\mathcal{E}$ as a function of temporality $\mathcal{T}$ and group size $N_{s}$.
@@
 Regardless of group size, the control energy reaches its minimum at $\mathcal{T}=1$, highlighting the optimality of ISA for herding noisy swarms along predefined paths.
-Figure 6: Indecisive Swarm Algorithm: a A schematic illustrates how temporality $\mathcal{T}$ influences a swarm of agents steered by a controller.
+The optimality of ISA can be attributed to its ability to combine the best features of ASA and LFSA.
+ASA agents quickly align with each other due to their averaging behavior, but they exhibit a strong order that requires the controller to apply high input intensities to reorient the agents in the preferred direction (Figure 7a) .
+For LFSA, due to the network hierarchy, the information from the controller can reach the follower nodes even if the input signal strength is low.
+However, fixed pairwise non-reciprocal interactions often cause the swarm to split into small clusters, preventing it from reaching the target (Figure 7b, SI video 11).
+ISA reduces the likelihood of cluster formation through network restructuring while maintaining sufficient flexibility to avoid the strong order of ASA agents.
+As a result, ISA agents successfully reach the target with significantly lower $I_{min}(t)$ compared to ASA (Figure 7c).
+Figure 6: Indecisive Swarm Algorithm.
+a A schematic illustrates how temporality $\mathcal{T}$ influences a swarm of agents steered by a controller.
 Each snapshot depicts the agents’ orientation, the controller’s relative position, and their interactions.
@@
 Control energy $\mathcal{E}$ achieves a minimum at $\mathcal{T}=1$, demonstrating the optimality of ISA for herding noisy agents along predefined trajectories.
+Figure 7: Schematic network representation of ASA, LFSA and ISA.
+The red node (node 1) represents the controller.
+a In ASA, due to the network being fully connected, a high input strength is required from the controller to align the agents in the preferred orientation.
+b In LFSA, due to the fixed pairwise non-reciprocal interactions, the network often splits in small clusters making it impossible to control.
+c ISA reduces the likelihood of cluster formation through network restructuring while maintaining sufficient flexibility to avoid the strong order of ASA agents.
 Discussion
@@
 This raises a key question: can temporality still offer advantages in the absence of knowledge about future changes?
-We demonstrate that for a specific class of control problems – herding of indecisive collectives – temporality can significantly reduce control energy, even without prior knowledge of topology changes.
-Herding differs from standard control problems in complex networks in two fundamental ways.
-First, while traditional controllability in the context of complex networks involves steering the system from any initial state to any desired state within the state space [8, 53], herding focuses on guiding all agents (nodes) to a fixed consensus state along a predefined trajectory.
-Second, unlike standard control, where the controller can choose which nodes receive input [8], in herding, the controller has no choice over which node receives the control input.
+We demonstrate that for a specific class of control problems–herding–temporality can significantly reduce control energy, even without prior knowledge of topology changes.
+While traditional controllability in the context of complex networks involves the ability of the controller in steering the system from any initial state to any desired state within the state space [8, 62], herdability focuses on guiding all agents (nodes) to a fixed consensus state along a predefined trajectory [54].
 Our analysis reveals that indecisive collective–stochastic temporal networks with restructuring timescales equal to system dynamics timescales ($\mathcal{T}=1$)–are optimal for minimizing control energy.
@@
 Instead, they function as a unified entity, sharing cognitive resources to analyze and anticipate the sheep’s behavior in real time [34, 66, 35].
-Systematically studying these sophisticated interactions – spanning verbal, physical, and visual modalities – could inspire advances in swarms robotics.
-Embedding distributed cognitive mechanisms across diverse agents, including drones, quadruped robots, and human or AI operators, could enable more adaptive, robust, and noise-tolerant systems capable of handling internal and external uncertainties effectively [58, 59].
+Systematically studying these interactions, spanning verbal, physical, and visual modalities, could reveal more rich complexities hidden in these multi-species control dynamics, and offering insights into principles of decentralized and stochastic collective control.
 Acknowledgements